# Structure of the Cayley-Dickson algebras

### v1

@@ X $[o_{1},o_{4},o_{2}]=-2o_{124}$ $[o_{1},o_{2},o_{4}]=2o_{124}$ $[o_{2},o_{1},o_{4}]=-o_{124}$
$[c,bc,d]=c(bc)d-c(bcd)=-bccd-cbdc=bd-c(bd)c=0\text{,}$
@@
$[bc,b,d]=bcbd-bc(db)=-cbbd+bc(bc(cd))=cd+cd(bc)(bc)=0\text{,}$
@@
This is isomorphic to an algebra documented in [12] which is part of a series of six algebras that break the symmetry of octonions whereby subalgebras of exceptional Lie algebra G2 provide automorphisms for these algebras. These have properties similar to the split octonions seen later, so were called the pseudo-octonions, designated ${\mathbb{P}_{k}}$ where $k$ is the number of non-associative triads $k=4,8,10,12,14,16$. The algebra above is ${\mathbb{P}_{4}}$ and only two other pseudo-octonion algebras, ${\mathbb{P}_{12}}$ and ${\mathbb{P}_{14}}$, are subalgebras of ultracomplex numbers. The number of occurrences of each subalgebra for the same algebras as shown in Table 4 is provided in Table 5.
@@
Defining $S_{m}$ and $O_{m}$ as the number of total ${\mathbb{P}_{k}}$ and ${\mathbb{O}}$ copies in ${\mathbb{U}_{m}}$ where the number of generators is $m+3$ then $S_{m+1}=7\times(O_{m}+S_{m})$. For convenience, $m=0$ case is taken to be ${\mathbb{U}_{0}}={\mathbb{O}}$, so $m\geq 0$.
@@
Table 5 shows these numbers for cases $m\in{\mathbb{N}}_{0}^{7}$. The number of blades or pure basis elements is $N_{m}=2^{m+3}-1$ so that the number of ${\mathbb{H}}$ in ${\mathbb{U}_{m}}$ is given by the number of paired blades, $\binom{N_{m}}{2}$, as seen in the table. The total of the number of triads in ${\mathbb{U}_{m}}$ is $\binom{N_{m}}{3}$ which is observed from Table 5 to be
$\binom{N_{m}}{3}=O_{m}+S_{m}+\binom{N_{m}}{2}\text{.}$
It is also observed that each ${\mathbb{P}_{k}}$ is divisible by 7, hence $S_{m}$ is divisible by 7, and that $S_{m}=(S_{m-1}+O_{m-1})\times 7$. Hence
$\frac{S_{m}}{S_{m-1}}=\frac{\binom{N_{m}}{3}-\binom{N_{m}}{2}}{\binom{N_{m-1}}{3}-\binom{N_{m-1}}{2}}\text{,}$
so the ratio of increase of the pseudo-octonion algebras equates to the rate of increase of the non-associate triads. By induction the series proceeds past ${\mathbb{U}_{7}}$ with the same ratio. Thus, the increase from ${\mathbb{U}_{m}}$ to ${\mathbb{U}_{m+1}}$ is $S_{m+1}=7\times(O_{m}+S_{m})$ and the octonions fill in the difference, $O_{n+1}=\frac{S_{n+2}}{7}-S_{n+1}$. ∎
@@ 4 Zero Divisors
$ac+bd=0\text{ and }ad+bc=0\text{,}$ (18)
with the second equation being an identity thus reducing the definition to just $ac=-bd$. Further, this is equivalent to Type 3 associativity and means the number of zero divisors at level $n$, $Z_{n}$, is
$Z_{n+1}=\left(\binom{2^{n}-1}{3}-\binom{2^{n}-1}{2}\right)\times 84\text{.}$
@@
Note that the reversed product $(c+d)(a+b)$ generates $-ac-bd=0$ so provide the same zero divisor pair. Finally, substituting for $a$, $ad=cbdd/d^{2}=-bc$ so that $ad+bc=0$ is an identity.
Hence zero divisor pairs are separated into two types for A and B non-associative triads that are both Type 3 associative. Octonions can not have zero divisors because the non-associative triads are X non-associative. Swapping the last two elements of the associator in Table 1 swaps Type 1 associativity and Type 2 associativity and has no correspondence with Type 3 associativity other than changing the sign. Hence A and B zero divisors may overlap because swapping the terms within brackets produces the same zero divisor pair. Table 5 shows 12 A non associative types for ${\mathbb{P}_{4}}$, which corresponds to 12 zero divisor pairs. The other pseudo-octonion algebras have mixures of A and B types adding up to more than 12 but since all ordered triples are included in the list of 28 triads then some A and B types overlap as zero divisors to give 12 distinct zero divisor pairs. Theorem 8 defines the number of pseudo-octonion algebras at each level as $S_{m+1}=7\times(O_{m}+S_{m})$ where $O_{m}$ and $S_{m}$ are the number of octonion and pseudo-octonion algebras at level $n=m+3$, respectively. This leads to $Z_{n+1}=(O_{n}+S_{n})\times 7\times 12$. ∎
@@
Calling (18) the prime zero divisor pair then $(-d+b)(c+a)$, is called the dual of the prime. The terminology dual is appropriate for zero divisors. Also $(a^{\prime}+b)(c+\mathop{\rm abs}\nolimits(db))$, with $a^{\prime}=bc\mathop{\rm abs}\nolimits(db)$, is called the extended zero divisor pair, and $(-\mathop{\rm abs}\nolimits(db)+b)(c+a^{\prime})$, is called both extended and dual. These four cases are called modes and whether they are zero divisors determines another level of structure for the ultronions.
@@ XXX All modes transform like the dual mode to other X types
Type 1 associativity [$\text{Type\leavevmode\nobreak\ 1 associativity }[b,d,c]$ $\rightarrow[b,a,c]\text{ which is Type\leavevmode\nobreak\ 1 associativity,}$,
[…]
@@
[…]
@@
[…]
@@ XXX $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ XXX
The first three elements are recognised as the pure quaternions and the first seven match the octonions [2]. Cawagas derives the $84$ zero divisor pairs shown in [5] and these are provided in graded form in Table 8 as $b$ and $c$ columns followed by prime, dual, extended and its dual triads as $d$, $a$, $db$ and $ab$ columns, respectively.
@@ 21 $o_{23}$ $o_{123}$ $o_{24}$ $o_{124}$ $o_{34}$ $o_{134}$
This example produces the zero divisor pair $(o_{1}-o_{1234})(o_{2}+o_{34})$ that provides an example of non-alternate associativity, $[a,a,b]\neq 0$ even though $[a,b,a]=0$,
@@
Now all other $bc$ pairs that don’t include $o_{5}$ can be covered. This involves triads with overlapping AAA and BBA cycles but the AAA cases still have four modes and the BBA have prime and extended modes only. In Table 12 the $d$ column provides the AAA structure and the $d_{1}$ and $d_{2}$ columns the BBA structure. The 21 sets of 3-cycles in Table 12 thus represent $21\times 3\times 4=252$ zero divisor pairs in the $d$ column and the $d_{1}$ and $d_{2}$ columns provide $21\times 3\times 2=126$ solutions each which means the table provides $252+126\times 2=504$ zero divisor pairs.
@@ 5 Split Cayley-Dickson Algebras
Equation (4) can also be defined with a positive number of $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. Such unitary generators are represented as $u_{i}$ with $u_{i}^{2}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. This provides a mixed sign basis with elements $o_{\alpha}u_{\beta}$ with $\alpha,\beta\in{\mathrm{B}(n)}$ such that $\alpha\cap\beta=\emptyset$. Lemma 3 shows that squaring such elements is $-1$ for even length of $\beta$ and is $+1$ for an odd length. This is now applied to each level of the graded algebras.
@@
and is described by the recurrence relation that each position after the first is the sum of the two elements above. For graded algebras the triangle enumerates Cayley-Dickson algebras starting with the reals, ${\mathbb{R}}$. Each row after this is the statement that the number of terms of any $k$-grade is given by the number of terms at grade $k$ without the generator $o_{n}$ or $u_{n}$ plus those that include this generator from grade $k-1$. Pascal’s triangle is horizontally symmetric because for each term of grade $k$, the complimentary term with grade $n-k$ contains the same $k$ indices removed from the $n$ generator case. For $n$ odd there are an equal number of terms with $o_{n}$ or $u_{n}$ as without in the complimentary terms. For $n$ even the same applies apart from the middle grade $k=n/2$. But $\binom{n}{n/2}$ is always even for $n$ even and by the definition of combinations it distributes the $n$ generators equally into an even set, half of which contain $o_{n}$ or $u_{n}$ and the others exclude it. Of course, the left-hand side has grade $k=0$ that specifies the basis number 1 while the right-hand side, the compliment of 1, has $k=n$ and always contains $o_{n}$ or $u_{n}$.
@@
The pure trace of the multiplication table for the algebra with $n$ generators is the sum of the squares of the $12^{n}-$ 1 pure basis terms, which is $-(2^{n}-1)$ for Cayley-Dickson algebra at level $n$. Thus for octonions this is $-7$.
@@
For any blade $o_{\alpha}u_{\beta}$, $\alpha$ and $\beta$ from disjoint sets of ${\mathrm{B}(n)}$ then Lemma 3 shows that $(o_{\alpha}u_{\beta})^{2}=\pm 1$, $-1$ if $\beta$ has even length and $+1$ if $\beta$ has odd length. Starting at $n=1$ the Lorentz numbers, $\widetilde{\mathbb{C}}$, have one blade, $u_{1}$, so has pure trace one. Split-quaternions, $\widetilde{\mathbb{H}}$, can have pure basis, $\{o_{1},u_{2},o_{1}u_{2}\}$ or $\{u_{1},u_{2},u_{12}\}$, both of which satisfy the theorem. By induction, the basis at level $n-1$ is assumed to have pure trace one and either $o_{n}$ or $u_{n}$ will multiplying all basis elements. By the binomial theorem recurrence relation the existing pure basis is doubled by multiplying all existing basis elements by $o_{n}$ and adding the $o_{n}$ (multiplication by 1). Extending each $\alpha$ in this way does not change the parity of $\beta$ so this doubles the pure trace count and adds the -1 trace part for $o_{n}$, keeping the pure trace to one.
Now applying the same recurrence relation with $u_{n}$ swaps even and odd parities of each $\beta$ in the new multiplied terms when excluding the unit, 1. Hence a trace of one at level $n-1$ has terms added with opposite trace parity giving a total parity of zero. But the extra term, $u_{n}$, from multiplication by 1, regains the pure trace of one. ∎
@@
All Cayley-Dickson algebras with any positive number of $\epsilon_{i}=-1$ for $i\in{\mathbb{N}}_{1}^{n}$ are isomorphic to each other so may be represented as the algebra of unitary generators, $u_{1},u_{2},\cdots,u_{n}$, called the split algebra, either $\widetilde{\mathbb{C}}$, $\widetilde{\mathbb{H}}$, $\widetilde{\mathbb{O}}$ or ${\rm\widetilde{\mathbb{U}}_{m}}$, $m=n-3>0$.
@@
From Theorem 1, all blades $x,y$ anti-commute, $xy=-yx$, if $x\neq y$, and by Lemma 3 have squares $x^{2}=\pm 1$, positive for an odd number of unitary generators and negative otherwise. A basis of the algebra is $x=o_{\alpha}u_{\beta}$, such that $\alpha\cap\beta=\emptyset$, $\forall\alpha,\beta\in{\mathrm{B}(n)}$. By Theorem 12, any positive number of unitary generators produce a basis with pure trace one. Selecting a set of $n$ of $\alpha+\beta$ so that $(o_{\alpha}u_{\beta})^{2}=1$ but the set elements are not all from the split algebra at level $n-1$ then from Lemma 1 this set generates the whole of ${\mathrm{B}(n)}$ and the corresponding basis elements can be relabelled as new generators $u_{1}^{\prime},u_{2}^{\prime},\ldots,u_{n}^{\prime}$. ∎
@@
Split-octonions have idempotents of the form $(1+u_{\alpha})$ with nilpotents $(1+u_{\alpha})(1-u_{\alpha})$ for $\alpha\in{\mathbb{N}}_{1}^{3}\cup\{123\}$. These are derived from the four positive square blades of $\widetilde{\mathbb{O}}$, $u_{\alpha}^{2}=1$ with $u_{123}^{2}=1$. Each idempotent can be multiplied on the left by the three blades that do not contain $u_{\alpha}$, which form a left ideal and gives $16$ nilponents. Multiplying by different terms of the nilpotent gives another zero divisor. For example, $(1+u_{1})$ has $(-u_{2}+u_{12})(u_{3}+u_{13})=0$, which satisfies the definition for a zero divisor since $u_{12}u_{3}u_{13}=-u_{2}$. Hence ideals overlap with zero divisors for split algebras.

### v2

@@ X $o_{1},o_{4},o_{2}]=-2o_{124} [$$o_{1},o_{2},o_{4}]=2o_{124} [$$o_{2},o_{1},o_{4}]=-2o_{124}$
$c,bc,d]=c(bc)d-c(bcd)=-bccd-cbdc=bd-c(bd)c=0\text{,}$
@@
$bc,b,d]=bcbd-bc(db)=-cbbd+bc(bc(cd))=cd+cd(bc)(bc)=0\text{,}$
@@
Each of these are isomorphic to an algebra documented in [10] which is part of a series of six algebras that break the symmetry of octonions whereby subalgebras of exceptional Lie algebra G2 provide automorphisms for these algebras. These have properties similar to the split octonions seen later, so were called the pseudo-octonions, designated $\mathbb{P}_{k}$} where $k$ is the number of non-associative triads $k=4,8,10,12,14,16$. The first case, $\mathbb{P}_{4}$}, was discovered[7] and enumerated[9] by Cawagas. Only this algebra and two other pseudo-octonion algebras, $\mathbb{P}_{12}$} and $\mathbb{P}_{14}$}, are subalgebras of ultracomplex numbers, the other three are not included. The number of occurrences of each subalgebra for the same algebras as shown in Table 4 is provided in Table 5. Octonions have seven quaternion subalgebras, seen in LABEL:eqn:qa7 for $(o_{1},o_{2},o_{3})$, and the pseudo-octionions have seven octonion-like algebras, split or not. Each quaternion-like algebra overlaps three times with other octonion-like rings. For example, $(o_{1},o_{2},o_{12})$ occurs in $(o_{1},o_{2},o_{3})$, $(o_{1},o_{2},o_{4})$ and $(o_{1},o_{2},o_{34})$. Hence the number of associative sub-algebras in sedonions is $(7+8)\times 7/3=35$ and from the table is can be seen that the number of overlapping quaternion-like algebras for each octonion-like subalgebra in $\mathbb{U}_{k}$} is $2^{k+1}-1$.
@@
Definition.
Define the number of blades or pure basis elements in $\mathbb{U}_{m}$} as $N_{m}=2^{m+3}-1$. Also define $H_{m}$, $S_{m}$ and $O_{m}$ as the number of total $\mathbb{H}$}, $\mathbb{P}_{k}$} and $\mathbb{O}$} copies in $\mathbb{U}_{m}$}, respectively. Each $\mathbb{P}_{k}$} and $\mathbb{O}$} consists of 28 unique non-associative triads, unique due to Theorem 4 for $\mathbb{O}$}. Hence the total number of triads in $\mathbb{U}_{m}$}, which is the number of faces in the $(m+2)$-simplex, is
$\binom{N_{m}}{3}=H_{m}+28\times(O_{m}+S_{m})$
@@
The number of $\mathbb{P}_{k}$} subalgebras in $\mathbb{U}_{m}$} is $S_{m+1}=7\times(O_{m}+S_{m})$. For convenience, $m=0$ case is taken to be ${\mathbb{U}_{0}}={\mathbb{O}}$, so $m\geq 0$.
@@
Table 5 shows these numbers for cases $m\in{\mathbb{N}}_{0}^{7}$. It is also observed that the count for each $\mathbb{P}_{k}$} is divisible by 7, hence $S_{m}$ is divisible by 7, and that $S_{m}=(S_{m-1}+O_{m-1})\times 7$, for $m\in{\mathbb{N}}_{0}^{7}$. By induction, assume this is true in general and prove the $S_{m+1}$ case. Since $\binom{N_{m}}{3}-H_{m}=28\times(O_{m}+S_{m})$ then
$\frac{S_{m+1}}{S_{m}}=\frac{\binom{N_{m+1}}{3}-H_{m+1}}{\binom{N_{m}}{3}-H_{m}}\text{.}$
so the ratio of increase of the pseudo-octonion algebras equates to the rate of increase of the non-associate triads. By induction the series proceeds past $\mathbb{U}_{7}$} with the same ratio and the octonions fill in the difference, $O_{m+1}=\frac{S_{m+2}}{7}-S_{m+1}$. ∎
@@ 4 Zero Divisors
[…]
with the second equation being an identity thus reducing the definition to just ac=bdac=-bd. Further, this is equivalent to Type 3 associativity and means the number of zero divisors for $\mathbb{U}_{n}$}, $Z_{m}$, is
$Z_{m}=\frac{1}{16}(N_{m}-1)(N_{m}-3)(N_{m}-7)\text{.}$
@@
Note that the reversed product $(c+d)(a+b)$ generates acbd=0-ac-bd=0 so provide the same zero divisor pair. Finally, substituting for $a$, $ad=cbdd/d^{2}=-bc$ so that $ad+bc=0$ is an identity.
Hence zero divisor pairs are separated into two types for A and B non-associative triads that are both Type 3 associative. Octonions can not have zero divisors because the non-associative triads are X non-associative. Swapping the last two elements of the associator in Table 1 swaps Type 1 associativity and Type 2 associativity and has no correspondence with Type 3 associativity other than changing the sign. Hence A and B zero divisors may overlap because swapping the terms within brackets produces the same zero divisor pair. Table 5 shows 12 A non associative types for $\mathbb{P}_{4}$}, which corresponds to 12 zero divisor pairs. The other pseudo-octonion algebras have mixures of A and B types adding up to more than 12 but since all ordered triples are included in the list of 28 triads then some A and B types overlap as zero divisors to give 12 distinct zero divisor pairs. Theorem 8 defines the number of pseudo-octonion algebras at each level as $S_{m+1}=7\times(O_{m}+S_{m})=\frac{7}{28}\left(\binom{N_{m}}{3}-H_{m}\right)$ with $H_{m}$ quaternion-like or non-trivial associative algebras in $\mathbb{U}_{m}$}. The sequence for $H_{n}$ shown in Table 5 is provided by M4415 (A006095) in [24] as the Guassian binomial with coefficients $[n,2]$ for $q=2$. This is $a(n)=(2^{n}-1)(2^{n-1}-1)/3$, $n>2$. Substituting $2^{n}=N_{m}+1$ and $n=m+3$ gives, for $m\geq 0$,
$H_{m}=\frac{1}{6}(N_{m}^{2}-N_{m})\text{.}$
Hence
$S_{m}$ $=\frac{7}{28}\left(\binom{N_{m-1}}{3}-H_{m-1}\right)$
$=\frac{7}{168}\big(N_{m-1}(N_{m-1}-1)(N_{m-1}-2)-N_{m-1}^{2}+N_{m-1}\big)$
$=\frac{1}{24}(N_{m-1}^{3}-4N_{m-1}^{2}+3N_{m-1})$
$=\frac{1}{24}N_{m-1}(N_{m-1}-1)(N_{m-1}-3)\text{.}$
Substituting $N_{m-1}=(N_{m}-1)/2$, gives $S_{m}=\frac{1}{192}(N_{m}-1)(N_{m}-3)(N_{m}-7)$. Since each $\mathbb{P}_{k}$} algebra has 12 zero-divisors [10] then $Z_{m}=12S_{m}$. ∎
@@
Calling (19) the prime zero divisor pair then (d+b)(c+a)(-d+b)(c+a), is called the dual of the prime. The terminology dual is appropriate for zero divisors. Also $(a^{\prime}+b)(c+\mathop{\rm abs}\nolimits(db))$, with $a^{\prime}=bc\mathop{\rm abs}\nolimits(db)$, is called the extended zero divisor pair, and $(-\mathop{\rm abs}\nolimits(db)+b)(c+a^{\prime})$, is called both extended and dual. These four cases are called modes and whether they are zero divisors determines another level of structure for the ultronions.
@@ XXX All modes transform like the dual mode to other X types
$\text{Type\penalty 10000\ 1 associativity }[b,d,c]$ $\rightarrow[b,a,c]\text{ which is Type\penalty 10000\ 1 associativity,}$
$\text{Type\penalty 10000\ 2 associativity }[b,c,d]$ $\rightarrow[b,c,a]\text{ which is Type\penalty 10000\ 3 associativity (reversed), and}$
$\text{Type\penalty 10000\ 3 associativity }[c,b,d]$ $\rightarrow[c,b,a]\text{ which is Type\penalty 10000\ 2 associativity (reversed).}$
@@
$[b,d,c]$ $=bdc-b(dc)=bdc+dcb\text{,}$ (${\rm P}1_{1}$)
$[c,b,d]$ $=cbd-c(bd)=cbd+bdc\text{,}=-bcd-dbc$ (${\rm P}3_{1}$)
$[b,d,bc]$ $=bd(bc)-b(d(bc))=bd(bc)-bcdb\text{,}$ (${\rm P}1_{2}$)
$[bc,b,d]$ $=bcbd-bc(bd)=cd-bc(bd)\text{.}$ (${\rm P}3_{2}$)
@@
$[b,\mathop{\rm abs}\nolimits(db),c]$ $=\pm(b(db)c-b(dbc))=\pm(dc-bdcb)\text{,}$ (${\rm E}1_{1}$)
$[c,b,\mathop{\rm abs}\nolimits(db)]$ $=\pm(cb(db)-c(b(db)))=\pm(bc(bd)-cd)\text{,}$ (${\rm E}3_{1}$)
$[b,\mathop{\rm abs}\nolimits(db),bc]$ $=\pm(b(db)(bc)-b(db(bc)))=\pm(-bcd-bd(bc)b)\text{,}$ (${\rm E}1_{2}$)
$[bc,b,\mathop{\rm abs}\nolimits(db)]$ $=\pm(bcb(db)-bc(b(db)))=\pm(-dbc-bcd)\text{,}$ (${\rm E}3_{2}$)
$[c,bc,\mathop{\rm abs}\nolimits(db)]$ $=\pm(c(bc)(db)-c(bc(db)))=\pm(d-bc(bd)c)\text{,}$ (${\rm E}2_{3}$)
$[bc,c,\mathop{\rm abs}\nolimits(db)]$ $=\pm(bcc(db)-bc(c(db)))=\pm(-d+bc(dbc))\text{.}$ (${\rm E}3_{3}$)
The following equivalences can be identified. ${\rm E}1_{1}\equiv{\rm P}1_{1}b$, ${\rm E}3_{1}\equiv{\rm P}3_{2}$, ${\rm E}1_{2}\equiv{\rm P}1_{2}b$, ${\rm E}3_{2}\equiv{\rm P}3_{1}$, ${\rm E}2_{3}\equiv{\rm P}3_{2}c$ and ${\rm E}3_{3}\equiv{\rm P}3_{1}$. The later one is because if (${\rm P}3_{1}$) is zero then bc(dbc)=(bcd)(bc)=d(bc)(bc)=dbc(dbc)=(bcd)(bc)=-d(bc)(bc)=d so that (${\rm E}3_{3}$) is zero or both are non-zero. This is used in Table 7 to transform all silos to the extended mode equivalents and to show that this behaves as shown in Figures 2 where $\star$ means non-zero.
@@ XXX $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ XXX
The first three elements are recognised as the pure quaternions and the first seven match the octonions [4]. Cawagas derives the $84$ zero divisor pairs shown in [7] and these are provided in graded form in Table 8 as $b$ and $c$ columns followed by prime, dual, extended and its dual triads as $d$, $a$, dbdb and $ab$ columns, respectively.
@@ 21 $o_{23}$ $o_{123}$ $o_{24}$ $o_{124}$ $o_{34}$ $o_{134}$
This example produces the zero divisor pair $(o_{1}-o_{1234})(o_{2}+o_{34})$ that provides an example of non-alternate associativity, $a,a,b]\neq 0$ even though $[a,b,a]=0$,
@@
Now all other $bc$ pairs that don’t include $o_{5}$ can be covered. This involves triads with overlapping AAA and BBA cycles but the AAA cases still have four modes and the BBA have prime and extended modes only. Table 6 shows that for $\mathbb{P}_{12}$} and $\mathbb{P}_{14}$} the type B triads provide half of the 12 zero divisors for the algebra. In Table 12 the $d$ column provides the AAA structure and the $d_{1}$ and $d_{2}$ columns the BBA structure. The 21 sets of 3-cycles in Table 12 thus represent $21\times 3\times 4=252$ zero divisor pairs in the $d$ column and the $d_{1}$ and $d_{2}$ columns provide $21\times 3\times 2=126$ solutions each which means the table provides $252+126\times 2=504$ zero divisor pairs.
@@ 5 Split Cayley-Dickson Algebras
Equation (5) can also be defined with a positive number of $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. Such unitary generators are represented as $u_{i}$ with $u_{i}^{2}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. This provides a mixed sign basis with elements $o_{\alpha}u_{\beta}$ with $\alpha,\beta\in{\mathrm{B}(n)}$ such that $\alpha\cap\beta=\emptyset$. Lemma 3 shows that squaring such elements is 1-1 for even length of $\beta$ and is +1+1 for an odd length. This is now applied to each level of the graded algebras.
@@
and is described by the recurrence relation that each position after the first is the sum of the two elements above. For graded algebras the triangle enumerates Cayley-Dickson algebras starting with the reals, $\mathbb{R}$}. Each row after this is the statement that the number of terms of any $k$-grade is given by the number of terms at grade $k$ without the generator $o_{n}$ or $u_{n}$ plus those that include this generator from grade k1k-1. Pascal’s triangle is horizontally symmetric because for each term of grade $k$, the complimentary term with grade nkn-k contains the same $k$ indices removed from the $n$ generator case. For $n$ odd there are an equal number of terms with $o_{n}$ or $u_{n}$ as without in the complimentary terms. For $n$ even the same applies apart from the middle grade $k=n/2$. But $\binom{n}{n/2}$ is always even for $n$ even and by the definition of combinations it distributes the $n$ generators equally into an even set, half of which contain $o_{n}$ or $u_{n}$ and the others exclude it. Of course, the left-hand side has grade $k=0$ that specifies the basis number 1 while the right-hand side, the compliment of 1, has $k=n$ and always contains $o_{n}$ or $u_{n}$.
@@
The pure trace of the multiplication table for the algebra with $n$ generators is the sum of the squares of the $2^{n}-1$ pure basis terms, which is $(2^{n}-1)$ for Cayley-Dickson algebra at level $n$. Thus for octonions this is 7-7.
@@
For any blade $o_{\alpha}u_{\beta}$, $\alpha$ and $\beta$ from disjoint sets of $\mathrm{B}(n)$} then Lemma 3 shows that $(o_{\alpha}u_{\beta})^{2}=\pm 1$, 1-1 if $\beta$ has even length and +1+1 if $\beta$ has odd length. Starting at $n=1$ the Lorentz numbers, $\widetilde{\mathbb{C}}$, have one blade, $u_{1}$, so has pure trace one. Split quaternions, $\widetilde{\mathbb{H}}$, can have pure basis, $\{o_{1},u_{2},o_{1}u_{2}\}$ or $\{u_{1},u_{2},u_{12}\}$, both of which satisfy the theorem. By induction, the basis at level n1n-1 is assumed to have pure trace one and either $o_{n}$ or $u_{n}$ will multiplying all basis elements. By the binomial theorem recurrence relation the existing pure basis is doubled by multiplying all existing basis elements by $o_{n}$ and adding the $o_{n}$ (multiplication by 1). Extending each $\alpha$ in this way does not change the parity of $\beta$ so this doubles the pure trace count and adds the -1 trace part for $o_{n}$, keeping the pure trace to one.
Now applying the same recurrence relation with $u_{n}$ swaps even and odd parities of each $\beta$ in the new multiplied terms when excluding the unit, 1. Hence a trace of one at level n1n-1 has terms added with opposite trace parity giving a total parity of zero. But the extra term, $u_{n}$, from multiplication by 1, regains the pure trace of one. ∎
@@
All Cayley-Dickson algebras with any positive number of $\epsilon_{i}=-1$ for $i\in{\mathbb{N}}_{1}^{n}$ are isomorphic to each other so may be represented as the algebra of unitary generators, $u_{1},u_{2},\cdots,u_{n}$, called the split algebra, either $\widetilde{\mathbb{C}}$, $\widetilde{\mathbb{H}}$, $\widetilde{\mathbb{O}}$ or $\rm\widetilde{\mathbb{U}}_{m}$}, m=n3>0m=n-3>0.
@@
From Theorem 1, all blades $x,y$ anti-commute, xy=yxxy=-yx, if $x\neq y$, and by Lemma 3 have squares $x^{2}=\pm 1$, positive for an odd number of unitary generators and negative otherwise. A basis of the algebra is $x=o_{\alpha}u_{\beta}$, such that $\alpha\cap\beta=\emptyset$, $\forall\alpha,\beta\in{\mathrm{B}(n)}$. By Theorem 12, any positive number of unitary generators produce a basis with pure trace one. Selecting a set of $n$ of $\alpha+\beta$ so that $(o_{\alpha}u_{\beta})^{2}=1$ but the set elements are not all from the split algebra at level n1n-1 then from Lemma 1 this set generates the whole of $\mathrm{B}(n)$} and the corresponding basis elements can be relabelled as new generators $u_{1}^{\prime},u_{2}^{\prime},\ldots,u_{n}^{\prime}$. ∎
@@
Split octonions have idempotents of the form $(1+u_{\alpha})$ with nilpotents $(1+u_{\alpha})(1-u_{\alpha})$ for $\alpha\in{\mathbb{N}}_{1}^{3}\cup\{123\}$. These are derived from the four positive square blades of $\widetilde{\mathbb{O}}$, $u_{\alpha}^{2}=1$ with $u_{123}^{2}=1$. Each idempotent can be multiplied on the left by the three blades that do not contain $u_{\alpha}$, which form a left ideal and gives $16$ nilponents. Multiplying by different terms of the nilpotent gives another zero divisor. For example, $(1+u_{1})$ has $(-u_{2}+u_{12})(u_{3}+u_{13})=0$, which satisfies the definition for a zero divisor since $u_{12}u_{3}u_{13}=-u_{2}$. Hence ideals overlap with zero divisors for split algebras.